# LoRA Done RITE: Robust Invariant Transformation Equilibration for LoRA Optimization

### v1

@@ A.5 Proof of Theorem 3
We continue from the proof of Theorem 3 and utilize Lemma 1. We already bound the second term in Theorem 3, so we only need to bound the third term.
@@
$\sum_{t=1}^{T}\|\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{P}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{P}}_{{\bm{A}}_{t}}^{\top}\|\leq\sum_{t=1}^{T}\operatorname*{Tr}(\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{P}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{P}}_{{\bm{A}}_{t}}^{\top})\leq\operatorname*{Tr}(\bar{{\bm{X}}}_{{\bm{A}}_{T}}^{-1})=\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{T}}^{1/2}).$
Summing up the second and third term, we get
$(2\eta+\frac{1}{\eta}\mu D_{\bm{A}}^{2}D_{\bm{B}}^{2})\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{T}}^{1/2}).$
Choosing $\eta=(1/\sqrt{2})\mu^{1/2}D_{\bm{A}}D_{\bm{B}}$, we have
$2\sqrt{2}\mu^{1/2}D_{\bm{A}}D_{\bm{B}})\operatorname*{Tr}(\bar{{\bm{V}}}_{{\bm{A}}_{T}}^{1/2})=O(D_{\bm{A}}D_{\bm{B}}GT^{-1/2}),$
@@
A.6 Gemma2 Experiment Result
Table 8: Gemma2 experimental results on LLM benchmarking datasets.
Model Optimizer HellaSwag ArcChallenge GSM8K OpenBookQA Avg.
Gemma2-2B Adam 93.50 77.47 51.10 87.00 77.27
LoRA+ 93.65 77.39 51.40 86.40 77.21
ScaledAdam 93.61 77.30 50.34 85.60 76.71
Lamb 93.53 76.88 48.07 85.00 75.87
LoRA-RITE 93.71 77.65 52.24 87.40 77.75
We conduct experiments on Gemma2-2B. Table 8 shows the experimental results. Our proposed method has a slight advantage over the baselines. We suspect that LoRA finetuning cannot give much performance improvement to Gemma2 on these tasks, thus the performance difference between each methods is low.

### v2

@@ A.6 Proof of Theorem 4
We prove Theorem 4 by bounding the right-hand side of Lemma 1. Since the first term is a constant and we already bound the second term in Theorem 3, here we only need to bound the third term.
@@
$\sum_{t=1}^{T}\|\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{P}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{P}}_{{\bm{A}}_{t}}^{T}\|\leq\sum_{t=1}^{T}\operatorname*{Tr}(\bar{{\bm{X}}}_{{\bm{A}}_{t}}^{-1}-{\bm{P}}_{{\bm{A}}_{t}}\bar{{\bm{X}}}_{{\bm{A}}_{t-1}}^{-1}{\bm{P}}_{{\bm{A}}_{t}}^{T})\leq\operatorname*{Tr}(\bar{{\bm{X}}}_{{\bm{A}}_{T}}^{-1}+c\rho_{{\bm{A}}_{T}}^{1/2}{\bm{I}})$
$\leq$ $\operatorname*{Tr}((\sum_{i=1}^{T}\bar{\nabla}{\bm{A}}_{i}^{\top}\bar{\nabla}{\bm{A}}_{i}+\rho_{{\bm{A}}_{T}}{\bm{I}})^{1/2})+c\operatorname*{Tr}(\rho_{{\bm{A}}_{T}}^{1/2}{\bm{I}})\leq(\sqrt{2}+c)\operatorname*{Tr}((\sum_{i=1}^{T}\bar{\nabla}{\bm{A}}_{i}^{\top}\bar{\nabla}{\bm{A}}_{i})^{1/2}),$
where the last inequality comes from Lemma 3.
Summing up the bound for the second and the third term, we get
$(2\eta+\frac{\sqrt{2}+c}{\eta}\mu D_{\bm{A}}^{2}D_{\bm{B}}^{2})\operatorname*{Tr}((\sum_{i=1}^{T}\bar{\nabla}{\bm{A}}_{i}^{\top}\bar{\nabla}{\bm{A}}_{i})^{1/2})$
Choosing $\eta=\mu^{1/2}D_{\bm{A}}D_{\bm{B}}$, we have
$(2+\sqrt{2}+c)\mu^{1/2}D_{\bm{A}}D_{\bm{B}}\operatorname*{Tr}((\sum_{i=1}^{T}\bar{\nabla}{\bm{A}}_{i}^{\top}\bar{\nabla}{\bm{A}}_{i})^{1/2})=O(D_{\bm{A}}D_{\bm{B}}GT^{-1/2}),$
@@
A.7 Training Loss Curve Visualization
To cross-validate the effectiveness of LoRA-RITE, we plot the training loss curve of each method for the Super-Natural instruction dataset and the OpenBookQA dataset. Figure 2 shows that LoRA-RITE has the lowest training loss, which demonstrates the effectiveness of our method.
(a) Super-Natural instruction
(b) OpenBookQA
Figure 2: The training loss curve for the Super-Natural instruction dataset and the OpenBookQA dataset.
A.8 Update Magnitude Visualization
To visualize the update magnitude of the two LoRA factors, we plot the update norm divided by the weight norm, $\|\delta{\bm{A}}\|/\|{\bm{A}}\|$ and $\|\delta{\bm{B}}\|/\|{\bm{B}}\|$.
Figure 3 and Figure 4 show that for conventional optimizers, factor ${\bm{A}}$ barely changes, while LoRA-RITE is able to learn the factor ${\bm{A}}$ effectively. This demonstrates the importance of transformation invariance.
(a) Update magnitude of the ${\bm{A}}$ factor
(b) Update magnitude of the ${\bm{B}}$ factor
Figure 3: The update magnitude of ${\bm{A}}$ and ${\bm{B}}$ for the Super-Natural instruction dataset.
(a) Update magnitude of the ${\bm{A}}$ factor
(b) Update magnitude of the ${\bm{B}}$ factor
Figure 4: The update magnitude of ${\bm{A}}$ and ${\bm{B}}$ for the OpenBookQA dataset.
A.9 Ablation Study on Different Ranks
To study the effect of different LoRA ranks, we conduct additional ablation study on different datasets.
As we can see from Table 8, higher rank generally improves LoRA performance, approaching full fine-tuning. This explains why the performance gap between LoRA-RITE and other methods narrows at higher ranks, as they all converge towards the results of full fine-tuning.
Additionally, one can observe that LoRA has inherent regularization properties. As noted in previous research (Chen et al., 2022), this means that sometimes a lower rank can actually lead to better performance. This effect depends on factors like model generalization and training data size. This explains why LoRA-RITE achieves better performance at rank 4 instead of 16 and why Lamb achieves better performance at rank 16 than rank 64.
Table 8: Ablation study on different ranks on Gemma-2B on the LLM benchmarking datasets.
Optimizer Rank HellaSwag ArcChallenge GSM8K OpenBookQA Avg.
Adam $r=4$ 81.83 42.32 20.92 63.0 52.02
$r=16$ 83.76 45.31 24.26 64.0 54.33
$r=64$ 84.56 46.67 26.08 67.0 56.08
ScaledAdam $r=4$ 81.95 44.80 21.15 63.0 52.73
$r=16$ 83.52 45.22 23.96 64.8 54.38
$r=64$ 84.42 48.21 26.61 67.0 56.56
Lamb $r=4$ 86.01 46.67 25.25 67.8 56.43
$r=16$ 86.60 47.35 26.76 68.0 57.18
$r=64$ 87.83 47.53 29.04 62.8 56.80
LoRA-RITE $r=4$ 87.08 49.57 29.49 70.4 59.14
$r=16$ 87.28 49.06 30.10 68.8 58.81
$r=64$ 87.89 49.91 31.46 68.8 59.52
A.10 Best Learning Rate for Different Optimizers
In Table 9, we list the best learning rate for each optimizer on the LLM benchmarking datasets. We observe that LoRA-RITE and Lamb usually prefer a larger learning rate than the other baselines.
Table 9: Best Learning Rate for Different Optimizers on LLM benchmarking datasets.
Model Optimizer HellaSwag ArcChallenge GSM8K OpenBookQA
Gemma-2B Adam 1e-5 5e-5 1e-5 5e-5
LoRA+ 1e-5 5e-5 1e-5 5e-5
ScaledAdam 5e-5 5e-5 1e-5 2e-4
Shampoo 1e-5 5e-5 5e-5 5e-5
Lamb 5e-3 5e-3 5e-3 5e-3
LoRA-RITE 2e-4 1e-3 2e-4 2e-4